OpenStudy (anonymous):

A company makes compact discs (CDs). The total cost, P pounds, of making n compact discs is given by the formula P = a + bn , where a and b are constants. The cost of making 1000 compact discs is £58 000. The cost of making 2000 compact discs is £64 000. Calculate the values of a and b.

OpenStudy (anonymous):

P=a+bn 58000=a+b(1000) 64000=a+b(2000)